# SoloSpeech: Enhancing Intelligibility and Quality in Target Speech Extraction through a Cascaded Generative Pipeline

### v2

@@ A.2 Conditional information in TSE
Appendix B Diffusion probabilistic model
@@
$q\left(x_{1:T}\mid x_{0}\right):=\prod_{t=1}^{T}q\left(x_{t}\mid x_{t-1}\right)\$ (2)
@@
Appendix C Model configurations
C.1 Audio compressor
@@
The STFT and its inverse (iSTFT) utilize a window size of 640 and a hop size of 320, resulting in latent representations at 50 Hz. The dimensionality $D$ is set to 128. In the encoder, a Conv2D layer is first applied to the complex-valued spectral input, with a kernel size of $3\times 3$, a zero-padding size of $1\times 1$, and 128 channels, followed by Group Normalization. Three TF-GridNet blocks are then applied. The embedding dimension for each T-F unit is 128, the kernel size for Unfold and Deconv1D layers is 1, the stride size for Unfold and Deconv1D layers is 1, the number of hidden units in bidirectional LSTMs (BLSTMs) in each direction is 256, the number of output channels in the $1\times 1$ Conv2D layers to obtain query and key tensors in the self-attention module is 512, and the number of heads in self-attention is 4. After that, we reshape the features and use a Conv1D layer with a channel size of 128. The PReLU function is used for all activations. The decoder is a mirror of the encoder. The model has a total of 49.3 million parameters.
@@
C.2 Target extractor
We developed three different model sizes for the target extractor: Small, Base, and Large. Unless otherwise specified, all results presented in this paper use the Base model. The performance of these three models can be found in Table H.5. In the Transformer column, the numbers denote the number of layers the model dimension, the number of heads, and the multiples of hidden size.
Table 5: Details of the target extractor models.
Model Params (M) MACs (G/s) Learning rate Diffusion Transformer Condition Transformer
Small 50.6 1.2 $2\times 10^{-4}$ 12, 384, 6, 2 12, 192, 6, 2
Base 200.8 4.8 $1\times 10^{-4}$ 12, 768, 12, 4 12, 384, 6, 4
Large 474.1 11.3 $5\times 10^{-5}$ 16, 1024, 16, 4 16, 512, 8, 4
C.3 Corrector
Following [19], we employed the Noise Conditional Score Network (NCSN++) architecture as described in [20], which is based on a multi-resolution U-Net structure operating on complex-valued STFT. The model has a total of 65.6 million parameters. For detailed information about the network, please refer to [20].
Appendix D Subjective Evaluations
For real-world TSE evaluation, we followed the protocol from [59], assessing both separation accuracy and perceptual quality through a Mean Opinion Score (MOS) test conducted with 12 human raters recruited via Prolific [^9] https://www.prolific.com/. We randomly selected 100 test samples from the CHiME-5 and RealSEP datasets, ensuring that each sample was annotated by three raters. Each listener rated the separation quality of two isolated tracks individually on a 1–5 scale (1 = Bad, 5 = Excellent), using the original mixture as a reference. We visualize the annotation interfaces for TSE in Figure 6,
Figure 6: Annotation UI for TSE.
Appendix E Baselines
E.1 TSE baselines
  • TD-SpeakerBeam [49]: A time-domain discriminative model that investigates strategies for improving speaker discrimination capability. We utilized their open-source code, which is available at https://github.com/BUTSpeechFIT/speakerbeam.
@@ E.1 TSE baselines
  • X-TF-GridNet [16]: a time-frequency domain discriminative model using TF-GridNet as its backbone. It extracts speaker embeddings using a U2-Net style network. We used their open-source code, which is available at https://github.com/HaoFengyuan/X-TF-GridNet.
  • USEF-TSE [17]: A speaker embedding-free discriminative model. We implemented the time-frequency model, i.e., USEF-TFGridNet, as described in the paper. The open-source code is available at https://github.com/ZBang/USEF-TSE.
@@
  • DPM-TSE [38]: A diffusion probabilistic model operating on the mel-spectrogram. We utilized a pretrained WavLM-based speaker verification network as the speaker embedding network. The source code is available at https://github.com/haidog-yaqub/DPMTSE.
@@
  • SoloAudio [23]: A latent diffusion model with a time-domain audio VAE as the latent features. We utilized a pretrained WavLM-based speaker verification network as the speaker embedding network. We utilized their open-source code, which is available at https://github.com/WangHelin1997/SoloAudio.
E.2 SS baselines
Appendix F Speech Separation with SoloSpeech
Although SoloSpeech is trained to extract one speaker from the mixture audio, we trained a target speech removal model to remove the target speaker and extract the other speaker. This approach allows for the separation of both speakers’ voices. As illustrated in Figure 7, the diffusion transformer in the target speech removal model receives four inputs: the noised latent, the mixture latent, the reference latent, and the current timestep.
During training, the target latent is obtained from the excluded speaker’s speech, while the reference latent is derived from the target speaker’s speech. In inference, the reference latent is obtained from the output of the target extractor model.
We used AdamW optimizer with a learning rate of $1\times 10^{-4}$, a weight decay of $1\times 10^{-4}$, a batch size of 32, and for 100 epochs. The model was trained on one NVIDIA A100-80GB GPU for three days.
Figure 7: Overview of the target speech removal model.
The whole pipeline of two-speaker separation using SoloSpeech is illustrated in Figure 8. The target speech is first extracted and then removed, resulting in a target speech and a interference speech.
Another naive manner to achieve speech separation is to simply reuse the original TSE model with the other speaker’s cue to extract the remaining speech. However, we find in practice that this approach does not produce high-quality complementary separation. We tried such approach that yielded an SI-SNR of 12.35 dB, below the 13.92 dB achieved by our separation method (in Table 3).
Figure 8: Overview of two-speaker separation pipeline.
Appendix G Computational Cost Analysis
We compare the computational cost of SoloSpeech with the state-of-the-art discriminative model (USEF-TSE) and the state-of-the-art generative model (SoloAudio), as summarized in Table 6. Both SoloSpeech and SoloAudio require considerably less training time than USEF-TSE due to the latent-space processing, enabling more practical large-scale training. While the real-time factor (RTF) of SoloSpeech is slightly higher, it remains suitable for offline or server-side deployment.
Since the primary motivation of this work is to achieve consistent improvements in the quality and intelligibility of TSE, particularly in terms of generalization, we argue that a modest increase in inference time is a reasonable trade-off. Furthermore, SoloSpeech’s modular architecture (compressor, extractor, corrector) supports the substitution of each component with more lightweight alternatives, which we plan to explore in future work.
Table 6: Comparison of computational cost on a single NVIDIA A100-80GB GPU.
Method Params (M) MACs (G/s) Training-time (hour per epoch) Real-time Factor (RTF)
USEF-TSE 19.7 129.0 2.80 0.34
SoloAudio 249.1 354.0 0.92 0.66
SoloSpeech 315.7 418.5 1.05 0.74
Appendix H Ablation Studies
H.1 Comparison of audio compressors
In the audio compressors reconstruction experiments, we utilized a set of audio quality metrics: STFT distance, MEL distance (as implemented in [35] with default parameters), PESQ, ESTOI, and SI-SNR. As shown in Table 7, the T-F Audio VAE consistently outperforms the Stable Audio VAE across all metrics, demonstrating superior audio reconstruction quality.

### v3

@@ II-B Conditional information in TSE
III Diffusion probabilistic model
@@
$q\left(x_{1:T}\mid x_{0}\right):=\prod_{t=1}^{T}q\left(x_{t}\mid x_{t-1}\right)\\$ (1)
@@
Figure 1: Overall pipeline of SoloSpeech.
IV SoloSpeech
IV-A Overall pipeline
The audio compressor transforms the speech signals into compact latent representations: $\boldsymbol{z}_{m}\in\mathbb{R}^{D\times L_{m}}$ for the mixture speech and $\boldsymbol{z}_{c}\in\mathbb{R}^{D\times L_{c}}$ for the conditional speech (see appendix IV-B for details). Here, $D$ denotes the dimensionality of the latent features for each frame, while $L_{m}$ and $L_{c}$ represent the number of frames in the mixture and conditional latents, respectively. The target extractor (detailed in appendix IV-C) then predicts the latent features of the target speech, $\boldsymbol{z}_{y}\in\mathbb{R}^{D\times L_{m}}$, from $\boldsymbol{z}_{m}$ conditioned on $\boldsymbol{z}_{c}$. The estimated target speech signal, $\boldsymbol{x}_{r}\in\mathbb{R}^{1\times T_{m}}$, is reconstructed using the decoder of the audio compressor. Finally, we introduce a corrector to refine the target speech, resulting in the refined signal $\boldsymbol{\hat{y}}\in\mathbb{R}^{1\times T_{m}}$. See appendix IV-D for details.
IV-B Audio compressor
The audio compressor operates on raw waveforms, compressing them into manageable sequence lengths. Current advanced audio compressors utilize time-domain VAEs with several convolutional blocks [50, 51, 52]. Inspired by the success of T-F domain modeling for mixture speech signals [53, 54], we propose a novel T-F domain VAE as the audio compressor. As illustrated in Figure 2, the encoder first applies the Short-Time Fourier Transform (STFT) [55] to transform the input audio signal $\boldsymbol{x}\in\mathbb{R}^{1\times N}$ into a complex spectrum $\boldsymbol{S}\in\mathbb{R}^{2\times F\times\frac{N}{h}}$, where $N$ denotes the audio length, $h$ the hop size of STFT, and $F$ the number of frequency bins. The real and imaginary components are concatenated along the first dimension.
Our VAE adopts TF-GridNet [53] as its backbone network, which comprises several TF-GridNet blocks. The encoder outputs a latent representation of shape $\mathbb{R}^{1\times 2D\times\frac{N}{h}}$, which is then split into two parts: the mean $\mu_{x}\in\mathbb{R}^{1\times D\times\frac{N}{h}}$ and the variance $\sigma_{x}\in\mathbb{R}^{1\times D\times\frac{N}{h}}$. The latent representation is obtained by sampling from this distribution, denoted as $z\sim\mathcal{N}(\mu_{x},\sigma_{x})$ where $z\in\mathbb{R}^{1\times D\times\frac{N}{h}}$.
The decoder mirrors the encoder structure, reconstructing the audio waveform using the inverse STFT (iSTFT). The VAE model is trained in a generative and adversarial manner [51]. The loss functions include: (i) a perceptually weighted multi-resolution STFT reconstruction loss [56]; (ii) an adversarial loss term with feature matching, utilizing five convolutional discriminators as in Encodec [57]; and (iii) a KL divergence loss term. See appendix LABEL:ap41 for details.
Figure 2: The audio compressor architecture.
IV-C Target extractor
The target extractor operates on the latent representations, aiming to extract the target latents $\boldsymbol{z}_{y}$ from the mixture latents $\boldsymbol{z}_{m}$, guided by the conditional latents $\boldsymbol{z}_{c}$. As shown in Figure 3a, we use a modified diffusion scheduler and velocity prediction [48], which have been shown to improve the purity and overall performance of sound extraction [32]. Details of the diffusion process could be found in appendix III. The main network for the target extractor employs a diffusion transformer with long skip connections (uDiT) [25], a modified version of DiT [58], as detailed in Figure 3b. These skip connections bridge shallow and deep DiT blocks, creating shortcuts for low-level features and streamlining the training of the entire velocity-prediction network. The input feature is a concatenation of the current noisy latent $\boldsymbol{z}_{t}$ and the mixture latent $\boldsymbol{z}_{m}$ in the latent dimension ($D$). The time feature and the conditional feature are fused into each block.
Figure 3: Architectures of the target extractor (a), Diffusion Transformer backbone (b) and uDiT block (c).
Unlike previous works that introduce speaker embeddings as conditional information [32, 25], our method jointly trains a condition transformer to extract the conditional features from the conditional latents $\mathbf{z}_{c}$, which are then attended to by the diffusion transformer using a cross-attention mechanism. This approach offers the following advantages: (i) it eliminates the need for extra data or labels to train a speaker embedding network; (ii) it preserves the sequential information of the conditional features, such as local dynamics and temporal structures; and (iii) it prevents potential mismatching between VAE latents and other types of features representing the target speech.
The condition transformer employs a ViT-like backbone [59], consisting of several ViT blocks. Each ViT block contains two LayerNorm layers, a Multi-Head Self-Attention (MHSA) layer, and MLP layers with SiLU activations [60]. The condition transformer generates a sequence of conditional features, which are then attended to by the uDiT blocks. As shown in Figure 3c, each uDiT block includes an MHSA module, a Multi-Head Cross-Attention (MHCA) module, and an MLP module. The conditional features are fused by the MHCA module, and the embedded time features are integrated into the MHSA module and the MLP module using Adaptive LayerNorm (AdaLN) layers. SiLU serves as the activation function, and we incorporate rotary positional embeddings (RoPE) [61] for the position encoding in both MHCA and MHSA.
IV-D Corrector
Diffusion models have been shown to outperform discriminative models for non-additive corruption types or when evaluated under mismatched conditions [22]. However, they often yield worse results on reference-based metrics [37], particularly for latent diffusion models, as VAE reconstruction determines the upper bound of audio quality. In addition, diffusion models may produce vocalizing and breathing artifacts in adverse conditions [38].
Inspired by the success of incorporating generative models to refine signal quality [20], we propose a T-F domain diffusion model-based corrector to address these issues. The corrector is designed to: (i) reduce artifacts caused by the front-end models; (ii) improve the signal quality from the audio compressor’s output by incorporating signal quality-based training objective functions; and (iii) correct errors introduced by the target extractor, such as speaker obfuscation and mispronunciation, thereby enhancing intelligibility.
IV-D1 Fast-GeCo Corrector
Inspired by Fast-GeCo [20], we employ a single-step diffusion model that allows us to directly optimize signal quality using the Scale-Invariant Signal-to-Noise Ratio (SI-SNR) loss function [62, 63]. As illustrated in Figure 4a, Fast-GeCo initially trains a multi-step diffusion model, which takes a noised signal as input, conditioned on the reference signal, mixture signal, and timestep. The noised signal depends on both the timestep and the reference signal. Subsequently, Fast-GeCo accelerates the traditional multi-step diffusion process by distilling it into a single-step operation, thereby producing the refined signal $\hat{\boldsymbol{y}}\in\mathbb{R}^{1\times T_{m}}$. Given the ground-truth target signal $\boldsymbol{y}$, the signal-step corrector is optimized by:
IV-D2 SoloSpeech Corrector
To address the above issues, in SoloSpeech, the refined signal is derived from the mixture signal rather than the reference signal, enabling the corrector to function similarly to a TSE task. As illustrated in Figure 4b, the corrector receives a noised mixture signal as input. To encourage the corrector to leverage more information from the mixture signal and enhance the intelligibility of the refined signal, we propose a simple yet effective augmentation approach for the reference signal. During the training phase, a random continuous span of the reference signal is masked. We set the masking ratio as $\theta$. This forces the model to recover the target signal conditioned on a corrupted reference signal, thereby compelling it to utilize additional information from the mixture signal.
Figure 4: Diagrams of Fast-GeCo corrector (a) and SoloSpeech corrector (b).
V Experiments
We conducted both in-domain and out-of-domain TSE experiments to evaluate the effectiveness and robustness of SoloSpeech. To examine the impact of each component in SoloSpeech, we performed ablation studies on TSE results involving different audio compressors, speaker conditions, and correctors.
V-A Datasets
To evaluate the performance of TSE, we trained models on the Libri2Mix 16k Hz dataset [30], which is simulated using WHAM!! noise data [64] and Librispeech utterances [65]. The dataset comprises two training subsets with 212 hours (train-360) and 58 hours (train-100) of audio, respectively. We utilized the train-360 subset for training. The dev set, containing 11 hours of audio, was used for validation. We then evaluated our models on the test set, which includes 11 hours of audio. During inference, the cue speech is a different utterance from the target speech by the same speaker, following the same data preparation as in TD-SpeakerBeam [66]. This dataset is also utilized for two-speaker SS experiments. Although Libri2Mix includes different speakers and background noises in its training and test sets, the recording conditions and speaking styles remain highly similar. Following [20], we assessed the trained models on out-of-domain data. Utterances from the Wall Street Journal (WSJ) corpus [62] served as the speech source, while noise audio from WHAM!!, MUSAN [67], and DEMAND [68] were used as noise sources for simulation. Each of these test sets has a duration of 5 hours. Following the approach in [64], noise was introduced by sampling a random SNR value from a uniform distribution ranging from -6 to +3 dB. Additionally, we generated a minimum-length version of the simulated data by removing any leading and trailing noise and truncating it to match the length of the shorter of the two speakers’ utterances. During inference, we selected a random speech sample of the target speaker different from the target speech as the cue speech.
We also evaluated the performance on two real-world datasets: CHiME-5 [69] and RealSEP [70]. CHiME-5 consists of real dinner party conversations recorded using distant microphones, featuring up to four overlapping speakers, expressive speech, non-verbal vocalizations (e.g., laughter, sigh), and background sound effects. RealSEP covers ten diverse real-world acoustic scenes, including music, sound effects, strong reverberation, and moving sound sources. Both datasets are highly challenging and contain conditions that are significantly different from those in Libri2Mix, demonstrating the robustness of our method in previously unseen scenarios. For each dataset, we randomly selected 100 mixtures (10–20 seconds) and their corresponding 10-second cue audio samples.
V-B Evaluation metrics
To evaluate the performance of SoloSpeech, we conducted TSE experiments using both quality and intelligibility metrics. The perceptual metrics include: (i) Perceptual Evaluation of Speech Quality (PESQ) [71]; (ii) Extended Short-Time Objective Intelligibility (ESTOI) [72]; (iii) Scale-Invariant Signal-to-Noise Ratio (SI-SNR) [62]; and (iv) Deep Noise Suppression Mean Opinion Score (DNSMOS) [73]. The first three are intrusive speech perceptual quality metrics, while DNSMOS is a non-intrusive metric that estimates the overall quality of the audio and naturalness. For the intelligibility metrics, we performed ASR on the speech files using the Whisper large-v3-turbo model [^2] https://huggingface.co/openai/whisper-large-v3-turbo [74] and calculated the Word Error Rate (WER) using the jiwer toolkit [^3] https://github.com/jitsi/jiwer. In addition, we measured the cosine similarity between the estimated speech and the ground truth using a pretrained WavLM-based speaker verification network [^4] https://huggingface.co/microsoft/wavlm-base-plus-sv [75] to assess speaker similarity (SIM).
TABLE I: Target speech extraction results on the Libri2Mix dataset. $\star$ are the results of models reproduced by us. D/G indicates if the model is discriminative or generative. Bold for the best result and underline for the second-best result.
Method D/G Perceptual Quality Naturalness Intelligibility
$\star$ TD-SpeakerBeam [66] D $1.66$ $0.70$ $9.21$ $3.14$ $0.21$ $0.93$
DPCCN [11] D $1.74$ $0.73$ $9.30$ $3.58$ - -
$\star$ X-TF-GridNet [17] D $1.72$ $0.72$ $9.85$ $3.42$ $0.19$ $0.93$
SUPERB-TSE [42] D $1.54$ $0.67$ $8.17$ $3.19$ $0.24$ $0.93$
SSL-MHFA [43] D $1.76$ $\underline{0.74}$ $\underline{10.60}$ $3.22$ $\underline{0.17}$ $\underline{0.94}$
$\star$ USEF-TSE [18] D $\underline{1.82}$ $0.72$ $10.17$ $3.48$ $\underline{0.17}$ $\underline{0.94}$
DiscreteTSE [35] G $1.29$ $0.69$ 39.95-39.95 - - -
DDTSE [39] G $1.60$ $0.71$ $7.60$ $\underline{3.74}$ - -
$\star$ Diff-TSE [19] G $1.56$ $0.64$ $8.35$ $3.53$ $0.25$ $0.89$
$\star$ DPM-TSE [32] G $1.50$ $0.60$ 1.56-1.56 $3.20$ $0.40$ $0.89$
$\star$ SoloAudio [25] G $1.56$ $0.63$ $2.64$ $3.71$ $0.35$ $0.93$
SoloSpeech (ours) G $\boldsymbol{1.89}$ $\boldsymbol{0.78}$ $\boldsymbol{11.12}$ $\boldsymbol{3.76}$ $\boldsymbol{0.15}$ $\boldsymbol{0.96}$
For the real-world TSE evaluation, we followed the protocol from [76], assessing both separation accuracy and perceptual quality through a Mean Opinion Score (MOS) test conducted with 12 human raters recruited via Prolific [^5] https://www.prolific.com/. We randomly selected 100 test samples from the CHiME-5 and RealSEP datasets, ensuring that each sample was annotated by three raters. Each listener rated the separation quality of isolated tracks on a 1–5 scale (1 = Bad, 5 = Excellent), using the original mixture as a reference. We report the mean scores along with $95\%$ confidence intervals.
V-C Model configurations
V-C1 Audio Compressor
@@
For the audio compressor, the STFT and its inverse (iSTFT) utilize a window size of 640 and a hop size of 320, resulting in latent representations at 50 Hz. The dimensionality $D$ is set to 128. In the encoder, a Conv2D layer is first applied to the complex-valued spectral input, with a kernel size of $3\times 3$, a zero-padding size of $1\times 1$, and 128 channels, followed by Group Normalization. Three TF-GridNet blocks are then applied. The embedding dimension for each T-F unit is 128, the kernel size for Unfold and Deconv1D layers is 1, the stride size for Unfold and Deconv1D layers is 1, the number of hidden units in bidirectional LSTMs (BLSTMs) in each direction is 256, the number of output channels in the $1\times 1$ Conv2D layers to obtain query and key tensors in the self-attention module is 512, and the number of heads in self-attention is 4. After that, we reshape the features and use a Conv1D layer with a channel size of 128. The PReLU function is used for all activations. The decoder is a mirror of the encoder. The model has a total of 49.3 million parameters.
@@
During training, we clipped the audio into 100-frame chunks. The VAE was trained using the AdamW optimizer, with a batch size of 16, base learning rate of $1.5\times 10^{-4}$, while the discriminators used a learning rate of $3\times 10^{-4}$. The weight decay was set to $1\times 10^{-3}$, and training was conducted for 10 epochs. It took two days on one NVIDIA A100-80GB GPU.
V-C2 Target Extractor
For the target extractor, we configure the Diffusion Transformer with 12 layers, a model dimension of 768, 12 attention heads, and a hidden-size multiplier of 4, and the Condition Transformer with 12 layers, a model dimension of 384, 6 attention heads, and a hidden-size multiplier of 4. The model has a total of 200.8 million parameters and was trained with a learning rate of $1\times 10^{-4}$, a weight decay of $1\times 10^{-4}$, a batch size of 32, and for 200 epochs. The training and inference steps were set to 1000 and 50, respectively, with the variance ranging from $8.5\times 10^{-4}$ to $1.2\times 10^{-2}$. During training, we applied data augmentation by randomly clipping the mixture and cue speech to lengths ranging from 3 seconds up to their original durations. The model was trained on one NVIDIA A100-80GB GPU for six days.
During inference, we found that classifier-free guidance (CFG) [77] did not improve performance in TSE. Large guidance scales even tended to produce overly smoothed outputs, with reduced speech detail and degraded intelligibility. As a result, we chose to exclude CFG in our pipeline. Instead, following [78], we generate 8 candidates and select the one with the highest speaker similarity score [^6] https://huggingface.co/microsoft/wavlm-base-plus-sv.
V-C3 Corrector
TABLE II: Target speech extraction results on the out-of-domain data. We trained the models on the Libri2Mix dataset and then tested them on these datasets.
[…]
WHAM! USEF-TSE [18] $9.15$ $0.23$ $3.19$ -
SoloAudio [25] $2.52$ $0.36$ $3.45$ -
SoloSpeech (ours) $\boldsymbol{10.68}$ $\boldsymbol{0.18}$ $\boldsymbol{3.72}$ -
MUSAN USEF-TSE [18] $8.96$ $0.25$ $0.20$ -
SoloAudio [25] $2.05$ $0.38$ $3.38$ -
SoloSpeech (ours) $\boldsymbol{10.40}$ $\boldsymbol{0.20}$ $\boldsymbol{3.70}$ -
DEMAND USEF-TSE [18] $9.70$ $0.19$ $3.25$ -
SoloAudio [25] $3.30$ $0.34$ $3.50$ -
SoloSpeech (ours) $\boldsymbol{11.41}$ $\boldsymbol{0.17}$ $\boldsymbol{3.72}$ -
CHiME-5 (real-world) USEF-TSE [18] - - $2.75$ $2.10\pm 0.15$
SoloAudio [25] - - $2.90$ $2.26\pm 0.14$
SoloSpeech (ours) - - $\boldsymbol{3.38}$ $\boldsymbol{2.93\pm 0.18}$
RealSEP (real-world) USEF-TSE [18] - - $2.69$ $2.04\pm 0.15$
SoloAudio [25] - - $2.52$ $1.87\pm 0.20$
SoloSpeech (ours) - - $\boldsymbol{3.15}$ $\boldsymbol{2.70\pm 0.18}$
For the corrector, we employed the same network as Fast-GeCo [20], which is based on a multi-resolution U-Net structure operating on complex-valued STFT, and set the reverse starting point $\tau=0.5$. The STFT utilizes a window size of 510 and a hop size of 128. During training, we clipped the audio into 256-frame chunks. We used a learning rate of $1\times 10^{-4}$, a batch size of 16 and 50 epochs, which took three days on one NVIDIA A100-80GB GPU.
V-D Baselines
  • TD-SpeakerBeam [66]: A time-domain discriminative model that investigates strategies for improving speaker discrimination capability. We utilized their open-source code [^7] https://github.com/BUTSpeechFIT/speakerbeam.
@@ V-D Baselines
  • X-TF-GridNet [17]: a time-frequency domain discriminative model using TF-GridNet as its backbone. It extracts speaker embeddings using a U2-Net style network. We used their open-source code [^8] https://github.com/HaoFengyuan/X-TF-GridNet.
  • USEF-TSE [18]: A speaker embedding-free discriminative model. We implemented the time-frequency model, i.e., USEF-TFGridNet, as described in the work. We used their open-source code [^9] https://github.com/ZBang/USEF-TSE.
@@
  • DPM-TSE [32]: A diffusion probabilistic model operating on the mel-spectrogram. We utilized a pretrained WavLM-based speaker verification network as the speaker embedding network. We used their open-source code [^10] https://github.com/haidog-yaqub/DPMTSE.
@@
  • SoloAudio [25]: A latent diffusion model with a time-domain audio VAE as the latent features. We utilized a pretrained WavLM-based speaker verification network as the speaker embedding network. We utilized their open-source code [^11] https://github.com/WangHelin1997/SoloAudio.
V-E Comparisons with state-of-the-art methods
To evaluate the efficiency of the proposed approach, we compared SoloSpeech with existing methods. We reproduced the results using the authors’ official settings. If the authors employed the same dataset as us, we report their metric values as presented in their original papers.
V-E1 In-domain Results
We trained and tested the models on the Libri2Mix dataset. As shown in Table I, SoloSpeech consistently outperformed existing methods across all metrics, including perceptual quality, naturalness, and intelligibility, achieving a 0.95 dB gain in SI-SNR. Compared to discriminative methods, SoloSpeech achieves significant improvements in DNSMOS, demonstrating its powerful ability to produce natural and high-quality audio. Furthermore, when compared with other generative approaches, SoloSpeech significantly outperforms them in perceptual and intelligibility metrics, highlighting its advantages. Higher values of DNSMOS and lower WER also indicate that our method produces fewer artifacts for both listeners and ASR systems.
V-E2 Out-of-domain Results
We trained the models on the Libri2Mix dataset and tested them on three out-of-domain datasets without any fine-tuning. As shown in Table II, SoloSpeech consistently outperforms existing methods in generalizing to unseen conditions. Specifically, SoloSpeech achieves SI-SNR gains of 16.7%, 16.1%, and 17.6% over the state-of-the-art discriminative method USEF-TSE on the WHAM!, MUSAN, and DEMAND datasets, respectively.
TABLE III: Impact of the modules. Unless specific statement, corrector is not applied. $\ast$ used fixed speaker embedding extracted from a pretrained WavLM-based speaker verification model. $\bullet$ finetuned the speaker embedding. $\diamond$ used features from a WavLM model.
Module Method SISNR $\uparrow$ DNSMOS $\uparrow$ WER $\downarrow$
Compressor Stable Audio VAE [51] $9.37$ $3.69$ $0.20$
T-F Audio VAE (ours) $\boldsymbol{10.10}$ $\boldsymbol{3.76}$ $\boldsymbol{0.18}$
Speaker Condition $\ast$ Speaker embedding [75] $9.25$ $3.72$ $0.20$
$\bullet$ Speaker embedding [75] $8.77$ $3.65$ $0.22$
$\diamond$ SSL-MHFA [43] $8.85$ $3.70$ $0.22$
Speaker network [66] $8.32$ $3.69$ $0.26$
Latent-space fusion (ours) $\boldsymbol{10.10}$ $\boldsymbol{3.76}$ $\boldsymbol{0.18}$
Corrector USEF-TSE [18] $10.17$ $3.48$ $0.17$
USEF-TSE w/ SoloSpeech corrector $11.06$ $3.55$ $0.16$
SoloAudio [25] $2.64$ $3.71$ $0.35$
SoloAudio w/ SoloSpeech corrector $6.58$ $3.72$ $0.22$
SoloSpeech w/o corrector $10.10$ $\boldsymbol{3.76}$ $0.18$
SoloSpeech w/ Fast-GeCo corrector [20] $10.55$ $\boldsymbol{3.76}$ $0.17$
SoloSpeech (ours) $\boldsymbol{11.12}$ $\boldsymbol{3.76}$ $\boldsymbol{0.15}$
V-E3 Real-world Results
We trained models on the Libri2Mix and tested them on two real-world datasets without any fine-tuning. As shown in Table II, SoloSpeech significantly outperforms baseline methods in generalizing to real-world challenging scenarios. Remarkably, SoloSpeech is able to generalize to more challenging unseen conditions. For example, handling more overlapping speakers (up to 4 in CHiME-5 vs. 2 in Libri2Mix) and much longer audio durations (up to 20 seconds). Moreover, it demonstrates strong robustness to factors such as expressive speech, non-verbal vocalizations, reverberation, and background sound effects.